# The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections

### v1

@@ 2.4.8 Pair production
At VHE, pair production $\gamma\gamma\xrightarrow{}e^{+}e^{-}$ can start to contribute to the SSC flux depletion. We account for this mechanism by a simplified treatment, assuming an isotropic distribution of the low-energy seed photons, as for the SSC calculation, and approximating the cross-section for pair production by a Dirac function at twice the threshold of the interaction. Then, VHE photons of frequency $\nu$ can produce pairs by interacting with low-energy photons at frequency $\nu_{\mathrm{seed}}=2\left(m_{\mathrm{e}}c^{2}\right)^{2}/\left(h^{2}\nu\right)$ and the characteristic timescale of this interaction is given by
@@
For very high frequencies $\nu$, the seed photons for pair production have a low frequency and $p_{\nu_{\mathrm{seed}}}^{\mathrm{syn}}+p_{\nu_{\mathrm{seed}}}^{\mathrm{SSC}}\simeq p_{\nu_{\mathrm{seed}}}^{\mathrm{syn}}$. The total emitted power per electron $p_{\nu}=p_{\nu}^{\mathrm{syn}}+p_{\nu}^{\mathrm{SSC}}$ is then corrected by a factor $\frac{t_{\gamma\gamma}(\nu)}{t_{\mathrm{dyn}}}\left(1-e^{-t_{\mathrm{dyn}}/t_{\gamma\gamma}(\nu)}\right)$, which in practice only attenuates the SSC component at high frequency where $t_{\gamma\gamma}(\nu)\ll t_{\mathrm{dyn}}$.
The implementation of the model allows to select the level of approximation for the SSC emission (synchrotron only, Thomson or full calculation) and to include or not the attenuation due to the pair production.
@@ 2.5 Observed flux
Once the emissivity in the comoving frame is known (§ 2.4), the flux measured by a distant observer with a viewing angle $\theta_{\mathrm{v}}$ is computed by an integration over equal-arrival time surfaces, taking into account relativistic Doppler boosting and relativistic beaming, as well as the effect of cosmological redshift. This leads to the following expression of the observer frame flux density ($\mathrm{erg}\cdot\mathrm{s}^{-1}\cdot\mathrm{cm}^{-2}\cdot\mathrm{Hz}^{-1}$) measured at time $t_{\rm obs}^{z}$ and frequency $\nu_{\rm obs}^{z}$(Woods & Loeb, 1999):
@@
where we use spherical coordinates $\left(r;\psi;\phi\right)$ with the polar axis equal to the line-of-sight, $\psi$ and $\phi$ the colatitude and longitude, and the jet axis in the direction $\left(\psi=\theta_{\mathrm{v}},\phi=0\right)$ (see Fig. 1), and where $z$ and $D_{\rm L}$ are the redshift and the luminosity distance of the source. Note that the angle $\psi$ differs from $\theta$ used in § 2.2 which is measured from the jet axis. The time $t$ (source frame) and frequency $\nu^{\prime}$ (comoving frame) are given by
@@
$\mathcal{D}(r,\psi,t)=\frac{1}{\Gamma(r,\psi,t)\left(1-\beta(r,\psi,t)\cos{\psi}\right)}$ (91)
@@
In practice, following the discretization of the jet structure described in § 2.1, the observer frame flux density $F_{\nu_{\rm obs}^{z}}\left(t_{\rm obs}^{z}\right)$ is computed as $F_{\nu_{\rm obs}^{z}}\left(t_{\rm obs}^{z}\right)=\sum_{i=0}^{N}F_{\nu_{\rm obs}^{z}}^{(i)}\left(t_{\rm obs}^{z}\right)$, where $F_{\nu_{\rm obs}^{z}}^{(i)}\left(t_{\rm obs}^{z}\right)$ are the contributions of the core jet ($i=0$) and lateral rings ($i=1\to N$). To compute each contribution, the thin shell approximation allows to reduce the double integral over $r$ and $\psi$ in Equation (2.5) to a simple integral over $r$, or equivalently over $\psi$ or $t$. In addition, to improve the computation time, the remaining integral on $\phi$ is computed analytically. This leads to
@@
$N_{\mathrm{e}}(r)=\zeta M_{\mathrm{ext}}(r)/m_{\mathrm{p}}$ is the number of shock-accelerated electrons per unit solid angle, and $p^{\prime(i)}_{\nu^{\prime}}(t)$ is the power per unit frequency and per electron in the comoving frame computed in § 2.4 and evaluated at time $t$ and comoving frequency $\nu^{\prime}=(1+z)\nu^{z}_{\mathrm{obs}}/\mathcal{D}_{i}\left(t^{z}_{\mathrm{obs}};t\right)$.
@@
and $\psi_{\mathrm{max},i}=\theta_{\mathrm{v}}+\theta_{\mathrm{max},i}$, i.e.
@@
Finally, we use the exact analytical calculation of the geometrical term $\Delta\phi_{i}$ defined by $\Delta\phi_{i}(\theta_{\mathrm{v}};\psi)=\int_{0}^{2\pi}\mathrm{d}\phi f_{i}(\theta_{\mathrm{v}};\psi;\phi)$, where $f_{i}(\theta_{\mathrm{v}};\psi;\phi)=1$ if the direction $\left(\psi;\phi\right)$ is contained in the component $i$ (core jet or ring) and 00 otherwise.
@@ 3.1 Dynamics and synchrotron emission
To test our calculation of the dynamics of the forward external shock and the associated synchrotron emission, we compare our results with two other models of the afterglow of a structured jet: the model presented in Gill & Granot (2018) and afterglowpy, a public Python module to calculate GRB afterglow light curves and spectra, based on Ryan et al. (2020).
@@
The first comparison is straightforward: with our default values of the normalization coefficients $K_{\nu}$ and $K_{\mathrm{P}}$, the assumptions (dynamics, microphysics, radiation, calculation of observed quantities) of our model in the pure synchrotron case are exactly the same as in Gill & Granot (2018). We checked that we reproduce exactly the different cases in their Fig. 4.
@@
The case of afterglowpy requires a more detailed comparison as there are differences in some treatments of the afterglow physics:
@@
(1) Early dynamics: in afterglowpy, even the early-time dynamics is computed assuming the self-similar Blandford & McKee regime, i.e. $\Gamma(t;\theta)\propto E(\theta)^{1/2}n_{0}^{1/2}t^{-3/2}$ (see § 2.1 in Ryan et al. 2020), whereas we include the coasting phase with $\Gamma\sim\mathrm{cst}$ before a smooth transition towards the self-similar regime at the deceleration radius, as described in § 2.2. Fig. 2 compares for a typical set of parameters the afterglow light curves from a top-hat jet viewed on-axis in radio, optical and X-rays obtained with our model (solid line), afterglowpy (dashed line) and our model where the dynamics has been forced to be in the self-similar regime at all times (dotted line). This allows to check that both models with the same self-similar dynamics agree well, except for the flux normalization as discussed below, and that the model with the full dynamics converge progressively towards the same solution: following the peak at the deceleration radius (at $\sim 10^{-2}\,\mathrm{days}$ in this example), the light curve obtained with the full dynamics smoothly converges towards the self-similar solution, and both light curves are identical at late times (typically after the jet break at $\sim 2\leavevmode\nobreak\ \mathrm{days}$ in this example). On the other hand, the self-similar approximation strongly over-estimates the fluxes until the deceleration radius and the full dynamics should always be included when considering early observations (see also Wang et al. 2024). In the case of GW 170817 discussed in the next section, the earliest detection by Chandra was obtained $\sim 9$ days after the merger, by which time the core jet dynamics had reached the self-similar regime.
@@
Figure 2: Synthetic afterglow light curves of a top-hat jet viewed on-axis, in the synchrotron-only radiation regime. Solid lines show the results obtained using our model with full dynamics treatment (using Equation (14)), while dotted lines follow the self-similar solution at all times. Dashed lines are obtained with afterglowpy without lateral expansion. The fluxes are computed in radio at $3\leavevmode\nobreak\ \mathrm{GHz}$ (black), in optical in r-band at $5.06\times 10^{14}\leavevmode\nobreak\ \mathrm{Hz}$ (blue) and in X-rays at $1\leavevmode\nobreak\ \mathrm{keV}$ (red). The parameters used for this figure are $\mathrm{E_{\mathrm{0,iso}}}=10^{52}\leavevmode\nobreak\ \mathrm{erg}$, $\theta_{\mathrm{c}}=4\leavevmode\nobreak\ \mathrm{deg}$, $n_{\mathrm{ext}}=10^{-3}\leavevmode\nobreak\ \mathrm{cm}^{-3}$, $\epsilon_{\mathrm{e}}=10^{-1}$, $\epsilon_{\mathrm{B}}=10^{-1}$, $p=2.2$, $D_{\mathrm{L}}=100\leavevmode\nobreak\ \mathrm{Mpc}$.
@@
(2) Late dynamics: contrary to afterglowpy, the present version of our model does not include lateral expansion of the ejecta. When comparing with afterglowpy, we therefore deactivate this option in the latter, which leads to the excellent late-time agreement seen in Fig. 2. The impact of the lateral spreading is expected to be very limited as long as the core jet is relativistic (see e.g. Woods & Loeb, 1999; Granot & Piran, 2012; van Eerten & MacFadyen, 2012; Duffell & Laskar, 2018), so that we stop the jet propagation when the core Lorentz factor reaches $\Gamma=2$ in the simulations of GW 170817’s afterglow discussed in the next section. In addition, we exclude observations after 400 days from the data set used for the afterglow fitting, which corresponds to $\Gamma\gtrsim 3-4$ for the core jet in our best fit models.
@@
(3) Flux normalization: afterglowpy relies on a scaling to boxfit (van Eerten et al., 2012), while in our model, this normalization is derived from analytical approximations detailed in Section 2. This leads to the difference of normalization seen in Fig. 2 when comparing afterglowpy with our model where the same self-similar dynamics is forced. The flux ratio varies between $1$ and $5$ at all wavelengths. When exploring a broad parameter space, varying the angle, energy injection, microphysical parameters, we checked that the flux ratio never exceeds $5$.
@@
To conclude this comparison, we note that afterglowpy and our model (in pure synchrotron mode) have been used by Kann et al. (2023) for two independent Bayesian inferences of the parameters of the afterglow of GRB 221009A using the same set of early observational data and converged towards very similar solutions.
@@
Figure 3: Spectral fits to the afterglow observations of GRB 190114C at $90\leavevmode\nobreak\ \mathrm{s}$ (black data points taken from MAGIC Collaboration et al. 2019b). Each color represents the best-fit spectrum for each of the possible radiative regimes (see text and Fig. 4). The spectral cases are detailed in Appendix B. This figure can be compared to Fig. 6 in Yamasaki & Piran (2022) showing very similar fits.
@@
Figure 4: Marginalized posterior distributions for the spectral fit of the afterglow observations of GRB 190114C at $t_{\mathrm{obs}}=90\leavevmode\nobreak\ \mathrm{s}$. The total (black) is subdivided between the slow (blue) and fast (red) cooling cases. The first row shows the four free parameters explored in this MCMC analysis ($\Gamma$, $\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}$, $B$ and $\gamma_{\mathrm{m}}$). On the second row, we show some derived quantities: $\widehat{\gamma}_{\mathrm{m}}/\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}/\gamma_{\mathrm{m}}$ and $B/\Gamma^{3}$. The bottom right panel shows the distribution of the radiative regimes found in the posterior sample (see Appendix B for their description). The best fit model for each radiative regime is shown in Fig. 3.
@@ 3.2 Synchrotron self-Compton emission
We now test the predictions of our model in the SSC regime. Several GRBs have been observed at VHE. Two events were particularly well followed-up at all wavelengths, GRB 190114C (MAGIC Collaboration et al., 2019a, b) and GRB 221009A (LHAASO Collaboration, 2023). In both cases, the VHE detection occurs at very early times but the analysis favors a dominant external forward shock origin for the emission. However, the case of GRB 221009A, which is an extreme GRB (the brightest ever detected by far) is quite complex. According to several studies, the early VHE emission and the later observations at lower frequencies probe different regions of a structured jet seen on-axis and may also include a significant contribution of the reverse shock at some frequencies (see e.g. Laskar et al., 2023; O’Connor et al., 2023; Gill & Granot, 2023; Sato et al., 2023; Zheng et al., 2024; Zhang et al., 2024; Derishev & Piran, 2024). We therefore decide to rather focus on GRB 190114C to test our model. This long GRB has a rich observational follow-up from radio to VHE at simultaneous times. Following MAGIC Collaboration et al. (2019a), we assume that the spectrum shows two distinct components.
@@
To test the radiative component of our model independently of the dynamics, we decide to analyse the observed spectrum at one fixed time, $t_{\mathrm{obs}}=90\leavevmode\nobreak\ \mathrm{s}$, following the approach proposed by Yamasaki & Piran (2022). Their spectral model is also based on Nakar et al. (2009), which allows an easy comparison. For this purpose, we use their values $K_{\mathrm{IC}}=1$ and $w_{\mathrm{KN}}=0.2$ in this subsection. We also consider a one-zone model of the emitting region, characterized by a radius $R$, a Lorentz factor $\Gamma$ and $N_{\mathrm{e}}$ accelerated electrons radiating in a magnetic field $B^{\prime}$. The electrons are injected with an initial distribution with a minimum Lorentz factor $\gamma_{\mathrm{m}}$ and a slope $p$. Following Yamasaki & Piran (2022), we fix $p=2.5$ and let four parameters free for the spectral fitting: $\Gamma$, $B^{\prime}$, $\gamma_{\mathrm{m}}$ and $\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}$. Then $R$ and $t^{\prime}_{\mathrm{dyn}}$ are deduced from $\Gamma$ and $t_{\mathrm{obs}}$, $n_{\mathrm{e}}^{\mathrm{acc}}$ from $B^{\prime}$, $\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}$, $\gamma_{\mathrm{m}}$ and $p$, and finally $N_{\mathrm{e}}$ from $n_{\mathrm{e}}^{\mathrm{acc}}$, $R$ and $\Gamma$ (see Section 2 in Yamasaki & Piran, 2022).
@@
We use the same observational data than Yamasaki & Piran (2022), shown in black in Fig. 3. We then fit the data using an MCMC procedure with log-uniform priors such that $3.51.3<\log\Gamma<$ 3.5; $40<\log\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}<$ 4; $-2<\log B<3$; $71<\log\gamma_{\mathrm{m}}<$ 7.
@@
The marginalized posterior distributions for these four parameters are provided on the first row of Fig. 4. As also discussed in Yamasaki & Piran (2022), several spectral solutions are possible. We show for further discussion the posterior distributions of $\widehat{\gamma}_{\mathrm{m}}/\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}/\gamma_{\mathrm{m}}$ and $B/\Gamma^{3}$, as well as the distribution of spectral cases in the bottom row of Fig. 4. The best fit model for each of the spectral cases is shown in Fig. 3. We find a high overall agreement with the results of Yamasaki & Piran (2022):
@@
  • the peak of the posterior distributions of $\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}$ and $\gamma_{\mathrm{m}}$ in Fig. 4 are very close to the assumed values in their Fig. 6 (as indicated by the vertical black dotted lines).
  • based on the analytical estimates from Derishev & Piran (2021), Yamasaki & Piran (2022) assume a strong correlation between the Lorentz factor and the magnetic field with $B/\Gamma^{3}\sim 10^{-7}$. Our MCMC exploration confirms this result but gives an estimate of the dispersion (see the distribution of $B/\Gamma^{3}$ in Fig. 4).
A detailed comparison shows that our results differ slightly on the exact parameter values, owing to some differing normalization constants. It should therefore be kept in mind that the absolute parameter values from spectral fits from different GRB afterglow studies should only be compared if the models used are the same.
@@
4 The VHE afterglow of GW 170817
Having validated the calculation of the dynamics and synchrotron + SSC emission in our model, we now present the results obtained by fitting the observations of the GRB afterglow of GW 170817, using Bayesian analysis.

### v2

@@ 2.4.8 Pair production
At VHE, pair production $\gamma\gamma\xrightarrow{}e^{+}e^{-}$ can start to contribute to the SSC flux depletion. We account for this mechanism by a simplified treatment, assuming an isotropic distribution of the low-energy seed photons, as for the SSC calculation, and approximating the cross-section for pair production by a Dirac function at twice the threshold of the interaction. Then, VHE photons of frequency $\nu$ can produce pairs by interacting with low-energy photons at a frequency $\nu_{\mathrm{seed}}=2\left(m_{\mathrm{e}}c^{2}\right)^{2}/\left(h^{2}\nu\right)$, and the characteristic timescale of this interaction is given by
@@
For very high frequencies $\nu$, the seed photons for pair production have a low frequency and $p_{\nu_{\mathrm{seed}}}^{\mathrm{syn}}+p_{\nu_{\mathrm{seed}}}^{\mathrm{SSC}}\simeq p_{\nu_{\mathrm{seed}}}^{\mathrm{syn}}$. The total emitted power per electron $p_{\nu}=p_{\nu}^{\mathrm{syn}}+p_{\nu}^{\mathrm{SSC}}$ is then corrected for by a factor $\frac{t_{\gamma\gamma}(\nu)}{t_{\mathrm{dyn}}}\left(1-e^{-t_{\mathrm{dyn}}/t_{\gamma\gamma}(\nu)}\right)$, which in practice only attenuates the SSC component at high frequency, where $t_{\gamma\gamma}(\nu)\ll t_{\mathrm{dyn}}$.
The implementation of the model allows us to select the level of approximation for the SSC emission (synchrotron only, Thomson, or full calculation) and to include or exclude the attenuation due to the pair production.
@@ 2.5 Observed flux
When the emissivity in the comoving frame is known (Sect. 2.4), the flux measured by a distant observer with a viewing angle $\theta_{\mathrm{v}}$ is computed by an integration over equal-arrival time surfaces, taking into account relativistic Doppler boosting and relativistic beaming, as well as the effect of cosmological redshift. This leads to the following expression of the observer frame flux density ($\mathrm{erg}\cdot\mathrm{s}^{-1}\cdot\mathrm{cm}^{-2}\cdot\mathrm{Hz}^{-1}$) measured at time $t_{\rm obs}^{z}$ and frequency $\nu_{\rm obs}^{z}$(Woods & Loeb 1999):
@@
where we use spherical coordinates $\left(r;\psi;\phi\right)$ with the polar axis equal to the line of sight, $\psi$ and $\phi$ the colatitude and longitude, and the jet axis in the direction $\left(\psi=\theta_{\mathrm{v}},\phi=0\right)$ (see Fig. 1), and where $z$ and $D_{\rm L}$ are the redshift and the luminosity distance of the source. The angle $\psi$ differs from $\theta$ used in Sect. 2.2, which is measured from the jet axis. The time $t$ (source frame) and frequency $\nu^{\prime}$ (comoving frame) are given by
@@
$\mathcal{D}(r,\psi,t)=\frac{1}{\Gamma(r,\psi,t)\left(1-\beta(r,\psi,t)\cos{\psi}\right)},$ (91)
@@
In practice, following the discretisation of the jet structure described in Sect. 2.1, the observer-frame flux density $F_{\nu_{\rm obs}^{z}}\left(t_{\rm obs}^{z}\right)$ is computed as $F_{\nu_{\rm obs}^{z}}\left(t_{\rm obs}^{z}\right)=\sum_{i=0}^{N}F_{\nu_{\rm obs}^{z}}^{(i)}\left(t_{\rm obs}^{z}\right)$, where $F_{\nu_{\rm obs}^{z}}^{(i)}\left(t_{\rm obs}^{z}\right)$ are the contributions of the core jet ($i=0$) and lateral rings ($i=1\to N$). To compute each contribution, the thin-shell approximation allows us to reduce the double integral over $r$ and $\psi$ in Eq. (2.5) to a simple integral over $r$, or equivalently, over $\psi$ or $t$. In addition, to improve the computation time, the remaining integral on $\phi$ is computed analytically. This leads to
@@
$N_{\mathrm{e}}(r)=\zeta M_{\mathrm{ext}}(r)/m_{\mathrm{p}}$ is the number of shock-accelerated electrons per unit solid angle, and $p^{\prime(i)}_{\nu^{\prime}}(t)$ is the power per unit frequency and per electron in the comoving frame computed in Sect. 2.4 and evaluated at time $t$ and comoving frequency $\nu^{\prime}=(1+z)\nu^{z}_{\mathrm{obs}}/\mathcal{D}_{i}\left(t^{z}_{\mathrm{obs}};t\right)$.
@@
and $\psi_{\mathrm{max},i}=\theta_{\mathrm{v}}+\theta_{\mathrm{max},i}$, that is,
@@
Finally, we use the exact analytical calculation of the geometrical term $\Delta\phi_{i}$ defined by $\Delta\phi_{i}(\theta_{\mathrm{v}};\psi)=\int_{0}^{2\pi}\mathrm{d}\phi f_{i}(\theta_{\mathrm{v}};\psi;\phi)$, where $f_{i}(\theta_{\mathrm{v}};\psi;\phi)=1$ when the direction $\left(\psi;\phi\right)$ is contained in the component $i$ (core jet or ring) and 00 otherwise.
@@ 3.1 Dynamics and synchrotron emission
To test our calculation of the dynamics of the forward external shock and the associated synchrotron emission, we compared our results with two other models of the afterglow of a structured jet: the model presented in Gill & Granot (2018), and afterglowpy, a public Python module for calculating GRB afterglow light curves and spectra, based on Ryan et al. (2020).
@@
The first comparison is straightforward: With our default values of the normalisation coefficients $K_{\nu}$ and $K_{\mathrm{P}}$, the assumptions (dynamics, microphysics, radiation, calculation of observed quantities) of our model in the purely synchrotron case are exactly the same as in Gill & Granot (2018). We verified that we exactly reproduced the different cases in their Fig. 4.
@@
The case of afterglowpy requires a more detailed comparison as there are differences in some treatments of the afterglow physics.
@@
(1) Early dynamics: In afterglowpy, even the early-time dynamics is computed assuming the self-similar Blandford & McKee regime, that is, $\Gamma(t;\theta)\propto E(\theta)^{1/2}n_{0}^{1/2}t^{-3/2}$ (see Sect. 2.1 in Ryan et al. 2020), whereas we included the coasting phase with $\Gamma\sim\mathrm{cst}$ before a smooth transition towards the self-similar regime at the deceleration radius, as described in Sect. 2.2. Fig. 2 compares for a typical set of parameters the afterglow light curves from a top-hat jet viewed on-axis in radio, optical, and X-rays obtained with our model (solid line), afterglowpy (dashed line), and our model where the dynamics was forced to be in the self-similar regime at all times (dotted line). This allows us to confirm that the two models with the same self-similar dynamics agree well, except for the flux normalisation, as discussed below, and that the model with the full dynamics progressively converges towards the same solution: Following the peak at the deceleration radius (at $\sim 10^{-2}\,\mathrm{days}$ in this example), the light curve obtained with the full dynamics smoothly converges towards the self-similar solution, and the two light curves are identical at late times (typically after the jet break at $\sim 2\leavevmode\nobreak\ \mathrm{days}$ in this example). On the other hand, the self-similar approximation strongly over-estimates the fluxes until the deceleration radius, and the full dynamics should always be included when considering early observations (see also Wang et al. 2024). In the case of GW 170817 discussed in the next section, the earliest detection by Chandra was obtained $\sim 9$ days after the merger, by which time, the core jet dynamics had reached the self-similar regime.
@@
Figure 2: Synthetic afterglow light curves of a top-hat jet viewed on-axis in the synchrotron-only radiation regime. The solid lines show the results obtained using our model with full dynamics treatment (using Eq. (14)). The dotted lines follow the self-similar solution at all times. The dashed lines are obtained with afterglowpy without lateral expansion. The fluxes are computed in radio at $3\leavevmode\nobreak\ \mathrm{GHz}$ (black), in optical in r band at $5.06\times 10^{14}\leavevmode\nobreak\ \mathrm{Hz}$ (blue), and in X-rays at $1\leavevmode\nobreak\ \mathrm{keV}$ (red). The parameters used for this figure are $\mathrm{E_{\mathrm{0,iso}}}=10^{52}\leavevmode\nobreak\ \mathrm{erg}$, $\theta_{\mathrm{c}}=4\leavevmode\nobreak\ \mathrm{deg}$, $n_{\mathrm{ext}}=10^{-3}\leavevmode\nobreak\ \mathrm{cm}^{-3}$, $\epsilon_{\mathrm{e}}=10^{-1}$, $\epsilon_{\mathrm{B}}=10^{-1}$, $p=2.2$, and $D_{\mathrm{L}}=100\leavevmode\nobreak\ \mathrm{Mpc}$.
@@
(2) Late dynamics: In contrast to afterglowpy, the current version of our model does not include lateral expansion of the ejecta. When we compared it with afterglowpy, we therefore deactivated this option in the latter, which led to the excellent late-time agreement shown in Fig. 2. The impact of the lateral spreading is expected to be very limited as long as the core jet is relativistic (see e.g. Woods & Loeb 1999; Granot & Piran 2012; van Eerten & MacFadyen 2012; Duffell & Laskar 2018), so that we stopped the jet propagation when the core Lorentz factor reached $\Gamma=2$ in the simulations of the afterglow of GW 170817 discussed in the next section. In addition, we excluded observations after 400 days from the data set used for the afterglow fitting, which corresponds to $\Gamma\gtrsim 3-4$ for the core jet in our best-fit models.
@@
(3) Flux normalisation: afterglowpy relies on a scaling to boxfit (van Eerten et al. 2012), while in our model, this normalisation is derived from analytical approximations detailed in Sect. 2. This leads to the difference in the normalisation shown in Fig. 2 when afterglowpy is compared with our model, where the same self-similar dynamics is forced. The flux ratio varies between $1$ and $5$ at all wavelengths. When exploring a broad parameter space, varying the angle, energy injection, and microphysical parameters, we verified that the flux ratio never exceeded $5$.
@@
To conclude this comparison, we note that afterglowpy and our model (in purely synchrotron mode) have been used by Kann et al. (2023) for two independent Bayesian inferences of the parameters of the afterglow of GRB 221009A using the same set of early observational data. The two codes converged towards very similar solutions.
@@
Figure 3: Spectral fits to the afterglow observations of GRB 190114C at $90\leavevmode\nobreak\ \mathrm{s}$ (black data points taken from MAGIC Collaboration et al. 2019b). Each colour represents the best-fit spectrum for each of the possible radiative regimes (see text and Fig. 4). The spectral cases are detailed in Appendix B. This figure can be compared to Fig. 6 in Yamasaki & Piran (2022), which shows very similar fits.
@@
Figure 4: Marginalised posterior distributions for the spectral fit of the afterglow observations of GRB 190114C at $t_{\mathrm{obs}}=90\leavevmode\nobreak\ \mathrm{s}$. The total (black) is subdivided between the slow (blue) and fast (red) cooling cases. The first row shows the four free parameters explored in this MCMC analysis ($\Gamma$, $\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}$, $B$ and $\gamma_{\mathrm{m}}$). In the second row, we show some derived quantities: $\widehat{\gamma}_{\mathrm{m}}/\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}/\gamma_{\mathrm{m}}$ and $B/\Gamma^{3}$. The bottom right panel shows the distribution of the radiative regimes found in the posterior sample (see Appendix B for their description). The best-fit model for each radiative regime is shown in Fig. 3.
@@ 3.2 Synchrotron self-Compton emission
We now test the predictions of our model in the SSC regime. Several GRBs have been observed at VHE. Two events were particularly well followed up at all wavelengths, GRB 190114C (MAGIC Collaboration et al. 2019a, b) and GRB 221009A (LHAASO Collaboration 2023). In both cases, the VHE detection occurred at very early times, but the analysis favoured a dominant external forward-shock origin for the emission. However, the case of GRB 221009A, which is an extreme GRB (the brightest ever detected by far) is quite complex. According to several studies, the early VHE emission and the later observations at lower frequencies probe different regions of a structured jet seen on-axis and may also include a significant contribution of the reverse shock at some frequencies (see e.g. Laskar et al. 2023; O’Connor et al. 2023; Gill & Granot 2023; Sato et al. 2023; Zheng et al. 2024; Zhang et al. 2024; Derishev & Piran 2024). We therefore decided to rather focus on GRB 190114C to test our model. This long GRB has rich observational follow-up from radio to VHE at simultaneous times. Following MAGIC Collaboration et al. (2019a), we assumed that the spectrum shows two distinct components.
@@
To test the radiative component of our model independently of the dynamics, we decided to analyse the observed spectrum at one fixed time, $t_{\mathrm{obs}}=90\leavevmode\nobreak\ \mathrm{s}$, following the approach proposed by Yamasaki & Piran (2022). Their spectral model was also based on Nakar et al. (2009), which allows an easy comparison. For this purpose, we used their values $K_{\mathrm{IC}}=1$ and $w_{\mathrm{KN}}=0.2$ in this subsection. We also considered a one-zone model of the emitting region, characterised by a radius $R$, a Lorentz factor $\Gamma$, and $N_{\mathrm{e}}$ accelerated electrons radiating in a magnetic field $B^{\prime}$. The electrons were injected with an initial distribution with a minimum Lorentz factor $\gamma_{\mathrm{m}}$ and a slope $p$. Following Yamasaki & Piran (2022), we fixed $p=2.5$ and let four parameters free for the spectral fitting: $\Gamma$, $B^{\prime}$, $\gamma_{\mathrm{m}}$, and $\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}$. Then, $R$ and $t^{\prime}_{\mathrm{dyn}}$ were deduced from $\Gamma$ and $t_{\mathrm{obs}}$, $n_{\mathrm{e}}^{\mathrm{acc}}$ from $B^{\prime}$, $\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}$, $\gamma_{\mathrm{m}}$, and $p$, and finally, $N_{\mathrm{e}}$ from $n_{\mathrm{e}}^{\mathrm{acc}}$, $R$, and $\Gamma$ (see Sect. 2 in Yamasaki & Piran 2022).
@@
We used the same observational data as Yamasaki & Piran (2022), shown in black in Fig. 3. We then fit the data using an MCMC procedure with log-uniform priors, such that $3.51.3<\log\Gamma<$ 3.5; $40<\log\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}<$ 4; $-2<\log B<3$; $71<\log\gamma_{\mathrm{m}}<$ 7.
@@
The marginalised posterior distributions for these four parameters are provided on the first row of Fig. 4. As also discussed in Yamasaki & Piran (2022), several spectral solutions are possible. We show for the further discussion the posterior distributions of $\widehat{\gamma}_{\mathrm{m}}/\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}/\gamma_{\mathrm{m}}$, and $B/\Gamma^{3}$, as well as the distribution of spectral cases in the bottom row of Fig. 4. The best-fit model for each of the spectral cases is shown in Fig. 3. We find a high overall agreement with the results of Yamasaki & Piran (2022).
@@
  • The peak of the posterior distributions of $\epsilon_{\mathrm{e}}/\epsilon_{\mathrm{B}}$ and $\gamma_{\mathrm{m}}$ in Fig. 4 is very close to the assumed values in their Fig. 6 (as indicated by the vertical dotted black lines).
  • Based on the analytical estimates from Derishev & Piran (2021), Yamasaki & Piran (2022) assumed a strong correlation between the Lorentz factor and the magnetic field with $B/\Gamma^{3}\sim 10^{-7}$. Our MCMC exploration confirms this result, but gives an estimate of the dispersion (see the distribution of $B/\Gamma^{3}$ in Fig. 4).
A detailed comparison shows that our results differ slightly for the exact parameter values, owing to some differing normalisation constants. It should therefore be kept in mind that the absolute parameter values from spectral fits from different GRB afterglow studies should only be compared when the models that were used are the same.
@@
4 Very high energy afterglow of GW 170817
Having validated the calculation of the dynamics and synchrotron + SSC emission in our model, we now present the results obtained by fitting the observations of the GRB afterglow of GW 170817 using a Bayesian analysis.